Operations Report – November 29th

rew 216 Operations Report 29-Nov-19
SOL: 5
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue, ongoing.
Notes on non-nominal systems:
Generator: run
Hours run: 12.5
From what time last night: 1758
To what time this morning: 0630
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 71 %
Diesel Reading – 39
Station Propane Reading – 57
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1462858
Water (static tank) (gal): 343
Static to Loft Pump used – yes
Water in Green Hab (gal): 267
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments: Weather day, no EVA ativity. HAB maintenance, mission planning for future EVAs and personal project/report time.
Summary of internet: No serious issues.
Summary of suits and radios: No issues.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: Routine watering.
Summary of Science Dome operations: Gypsum processing continued.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved.
Summary of any observatory issues: None, weathered out.
Summary of health and safety issues: HSO broke sim to address medical issue at Base Sopport.
Questions, concerns and requests to Mission Support: Need to replace 1 bottle of cough suppresent/expectorant consumed.

Copyright © The Mars Society. All rights reserved. | Main Site